OpenStudy (goldenfreddy):

Which decimal is equivalent to ? A. 5.45 B. 5.9 C. 5.92 D. 5.95

OpenStudy (goldenfreddy):

here is the rest of the question

OpenStudy (goldenfreddy):

@AlexandervonHumboldt2

OpenStudy (trojanpoem):

\[a\frac{ b }{c } = ac + b/c\] 5 * 20 + 9/20 = 109/20 109/ 10 * 2= 10.9/2 = 0.9/2 + 10/2 = 5.45